OpenStudy (anonymous):

A certain species of tree grows an average of 0.5 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 200 centimeters tall.

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

i got a(n) = 200+ 0.5n i wanna know if thats correct

OpenStudy (anonymous):

Un=200+(n-1)0.5 or Un=200+0.5n

OpenStudy (anonymous):

You are correct

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

But when we're talking about sequence its better to use Un and not a(n)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You're welcome